OpenStudy (anonymous):

(x+a)^2-(x+b)^2=a^2-b^2

OpenStudy (anonymous):

what do you want?????

OpenStudy (anonymous):

pls aswer equiation

OpenStudy (anonymous):

use the identity for Difference of Squares i.e. m^2 - n^2 = (m+n)(m-n)

OpenStudy (anonymous):

\[2ax+2bx=0 Use factor ??\]

OpenStudy (anonymous):

(x+a)^2-(x+b)^2=a^2-b^2 [(x+a)+(x+b)]*[(x+a)-(x+b)]= (a+b)*(a-b) [2x+a+b]*[a-b]=(a+b)*(a-b) divide both sides by (a-b) and u get 2x+a+b=a+b 2x = a+b-a-b 2x =0 x=0/2 x=0

OpenStudy (anonymous):

Thx so much

OpenStudy (anonymous):

You can see that if we put x=0 in the original equation, we get (0+a)^2-(0+b)^2=a^2-b^2 which gives a^2-b^2=a^2-b^2

OpenStudy (anonymous):

u r welcome ☺

OpenStudy (anonymous):

yea can see thx